OpenStudy (anonymous):

I am trying to find the volume of a quarter sphere. Can anyone tell me how to calculate this?

jimthompson5910 (jim_thompson5910):

volume of full sphere is \[\Large V = \frac{4}{3}\pi*r^3\]

jimthompson5910 (jim_thompson5910):

Multiply both sides by 1/4 \[\Large V = \frac{4}{3}\pi*r^3\] \[\Large \frac{1}{4}*V = \frac{1}{4}*\frac{4}{3}\pi*r^3\] \[\Large \frac{1}{4}*V = \frac{1}{\cancel{4}}*\frac{\cancel{4}}{3}\pi*r^3\] \[\Large \frac{1}{4}*V = \frac{1}{3}\pi*r^3\]

jimthompson5910 (jim_thompson5910):

So the volume of a quarter-sphere is \[\Large \frac{1}{3}\pi*r^3\]